OpenStudy (anonymous):

how do I change parametric equations to slope-intercept form? specific question below.

OpenStudy (anonymous):

x = -4t +3 y = 5t -3

DivineSolar (divinesolar):

Maybe this example will help? Easy example X=t-10 Y=t+2 Work X+10=t Y-2=t X+10=Y-2 Answer: X+12=Y

OpenStudy (jdoe0001):

in short, you use the "substitution" method for the parameter

DivineSolar (divinesolar):

1: X=5t-7 PAREMETRIC EQUATION Y=10t-15 Step 2: X+7=5t Y+15=10t STEP 2 and 3 (ISOLATING T) Step 3: (X+7)/5=t (Y+15)/10=t Step 4: (X+7)/5=(Y+15)/10 (SETTING EQUATIONS EQUAL TO EACH OTHER Step 5: ((X+7)/5)*10=Y +15 SOLVING FOR Y (10X+70)/5=Y+15 (2X+14)-15=Y 2X+1=Y MX+B=Y

DivineSolar (divinesolar):

This should help better.

OpenStudy (anonymous):

you just get them all set equal to t and then set those together?

DivineSolar (divinesolar):

Step 1 Parametric equation, 2- isolate the t, 3 same steps, 4 set the equations equal to each other, 5 solve for y

OpenStudy (anonymous):

thats a lot easier than I thought, thank you both!

DivineSolar (divinesolar):

Sure thing!

OpenStudy (jdoe0001):

\(\begin{array}{llll} x = -4t +3\implies x-3=-4t\implies \cfrac{x-3}{-4}={\color{brown}{ t}}\\ y = 5{\color{brown}{ t}} -3\implies y=5\left({\color{brown}{ \cfrac{x-3}{-4} }} \right)-3 \end{array}\)

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw